Find a polynomial function​ f(x) of least degree having only real coefficients and zeros as given. Assume multiplicity 1 unless otherwise stated.
zero of 3​ (multiplicity 2) and zero 6i.
f(x)=

Answer by ikleyn(52790)

.

    f(x) = (x-3)^2*(x-6i)*(x+6i) = (x-3)^2*(x^2+36).    ANSWER

Solved.

-----------------

The lesson to learn


        You should include complex conjugate root with each complex root.

        The rest is simple arithmetic.